A certain sample of gold ore is 6.50% gold and the rest is other stuff. How many grams of this gold ore would you need to refine in order to obtain 6.54x 10^46 atoms of gold
Answer #1

molar mass of Au = 197.0 g/mol

This is mass of 6.022*10^23 atoms

use:

mass of 1 atom = molar mass /(6.022*10^23)

mass of 1 atom = 197.0/(6.022*10^23)

mass of 1 atom = 3.271*10^-22 g

mass of 6.54*10^46 atom = mass of 1 atom * total number of atoms

mass of 6.54*10^46 atom = 3.271*10^-22 * 6.54*10^46

mass of 6.54*10^46 atom = 2.139*10^25 g

Now use:

Mass % of gold = mass of gold * 100 / mass of ore

6.50 = (2.139*10^25 g)*100 / mass of ore

mass of ore = 3.29*10^26 g

Answer: 3.29*10^26 g